OpenStudy (anonymous):

(-5,8) (2,-5) Find the distance between the pair of points with the given coordinates.

OpenStudy (rajee_sam):

do you know the distance between two points formula? Distance = \[\sqrt{(x _{2}-x _{1})^{2} + (y _{2}-y _{1})^{2}}\]

OpenStudy (anonymous):

so thats how i set it up an solve for it with that formula

OpenStudy (rajee_sam):

now consider one point as \[\left( x _{1} ,y _{1}\right) \] and the other point as \[\left( x _{2},y _{2} \right)\]

OpenStudy (rajee_sam):

yes

OpenStudy (anonymous):

\[2\sqrt{5}\]